OpenStudy (anonymous):

-3 over 8 - 1 over 10 A. 2 over 9 B. -19 over 40 C. -3 over 8 D, 19 over 40

OpenStudy (anonymous):

Giving medal to whoever can answer this correctly!

sam (.sam.):

Common denominator

OpenStudy (anonymous):

How do I figure that out?

sam (.sam.):

\[-\frac{3}{8} -\frac{1}{10} \\ \\ =-\frac{3}{8} (\frac{\times10}{\times 10})-\frac{1}{10} ( \frac{\times8}{\times8}) \\ \\ =\frac{-30-8}{80}\]

OpenStudy (anonymous):

o_O

OpenStudy (anonymous):

Im confused.:L

sam (.sam.):

Well you just multiply whatever the other fraction's denominator to get the fractions together \[-\frac{3}{\color{red}{8}} -\frac{1}{\color{blue}{10}}\] \[(-\frac{3}{8} \times \color{blue}{\frac{10}{10}}) -(\frac{1}{10}\color{red}{\times \frac{8}{8}})\] \[-\frac{30}{80} -\frac{8}{80}\] \[\frac{-30-8}{80}\] Then add -30 to -8

OpenStudy (anonymous):

I got -38 :S

OpenStudy (anonymous):

well positive 38

OpenStudy (anonymous):

IM so dumb>.>

sam (.sam.):

Negative, not positive

OpenStudy (anonymous):

Oh so positive. so would it be 3 over 8?

sam (.sam.):

Then find the common factor of -38 and 80, can you find it?

sam (.sam.):

nope,

OpenStudy (anonymous):

How would I find that?

sam (.sam.):

You gotta find the number that, when -38 and 80 is divided by that number, you get a whole number. Try dividing 2 for both -38 and 80

OpenStudy (anonymous):

Im not sure what you mean.:[

OpenStudy (anonymous):

2 for both -38 and 80?

sam (.sam.):

Yup

OpenStudy (anonymous):

What do you mean by that?

sam (.sam.):

-38 divide 2 you get -19 80 divide 2 you get 40

OpenStudy (anonymous):

Ohhh.

sam (.sam.):

Because you have to simplify it, by dividing 2 you're actually making the numbers smaller

OpenStudy (anonymous):

I see. can you chck the answer to this problem? I got it just want to make sure it's correct.

sam (.sam.):

ok

OpenStudy (anonymous):

1.4 - 5.5 - 1.73 + 1.8 - 1.09 -5.12 is my answer. you've got to simplify.

sam (.sam.):

That's correct

OpenStudy (anonymous):

Thank You!:p

sam (.sam.):

yw